Find the principal needed now to get the given amount, that is, find the present value.
To get $500 after 3 years at 5% compounded quarterly
The present value of $500 is $?

Answer by rfer(16322)

500=x(1.0125)^(4*3)
500=x(1.161)
500/1.161=x
x=$431